# Direct measurement of forces in air-based acoustic levitation systems

### v1

@@ I Introduction
where $r$ is the center-to-center distance and $E_{0}$ represents the acoustic energy density.Silva and Bruus (2014); Lim, VanSaders, and Jaeger (2024) Note that the magnitude of the secondary scattering force is proportional to the product of the two particles’ volumes. While attractive in the nodal plane, the interaction is repulsive normal to that plane, i.e. along the levitation axis. When many objects are levitated, this results in the formation of a close-packed, quasi-two-dimensional membrane or raft.Lim et al. (2022)
@@ II.3 Force detection, probe tips, and calibration
A convenient way to calibrate the forces experienced by the probes employs a $\upmu$g-resolution scale (Sartorius CPA26P). During calibration, the plastic handle at the far end of the cantilever wire is attached to a micrometer oriented along the vertical, $z$-direction and lowered so that it barely makes contact with the scale’s platform. The scale’s doors are closed as much as possible, and openings covered with plastic sheeting, to minimize noise from air currents in the room. Incrementally pressing the probe tip onto the scale, mass readings from the scale are recorded and converted to force. The force-deflection data are linear to excellent approximation and fit to a line. The resulting slope is the constant that relates maximum cantilever deflection (or deflection at the probe tip) to measured force. A sample calibration curve is plotted in Fig. 3(c).
@@ III Mapping an acoustic cavity
The results of a mapping of this kind are shown in Fig. 4. Here a probe shaped like that pictured in Fig. 4(a) was used, with a 900 $\upmu$m diam. sphere as the tip. The sphere was positioned with its equator in the nodal plane, keeping the rest of the probe tip out of plane. Using a larger probe has the advantage of experiencing a larger primary acoustic force (as primary forces scale with the volume of the scatterer Wu et al. (2023), see Eq. 1) and also makes it easier to locate in an image, which results in increased relative resolution for determining force values. This probe tip was moved from the center of the cavity along the $x$-axis to a distance greater than $\lambda/2$ and back twice, for four separate trials. For each measurement, we moved the probe into position and then waited for the acoustic pressure control to adjust to changed resonance conditions before recording observations. Videos were recorded with the field on and off for each position. For this and other experiments, image processing to locate probe positions in images was performed with Meta’s Segment Anything model.Kirillov et al. (2023)
@@ IV Measuring the scattering force
Figure 5: Measuring the acoustic scattering force between two objects in the acoustic trap. A T-shaped probe is moved towards a force sensor with a 180 $\upmu$m spherical tip and then pulled away. The spherical probe is on a cantilever to measure forces. The probe positions are tracked to experimentally determine the acoustic scattering force between the probes. (a) Images from the experiment, viewed from below. (a, top) Start of the experiment when the probes are well-separated. (a, bottom) Middle of the experiment when the two probes are almost in contact. Note that the spherical probe has moved slightly to the left as a result of attraction to the larger probe. (b) Schematic of the experimental procedure. The T-probe moves to approach the spherical probe; the spherical probe is attached to a spring-like cantilever to measure the attractive scattering force between the two objects. (c) Plot of force vs. distance for this experiment (red) and corresponding simulations (blue). Five trials of the experiment were performed, and the data binned by inter-probe distance (error bars: standard error). The scattering force decreases sharply as distance increases. The gray bar marks the noise floor for the experimental force sensor. The force is normalized to allow for comparison between experiment and simulation. The shape of the simulation results is similar to the experimental results. Two triangles are drawn to represent $r^{-3}$ (left) and $r^{-4}$ (right) power laws; these correspond to a simple analytical model of the scenario.
@@ VI Microscale mechanical testing of levitated granular materials
Tension force and displacement data for this experiment are shown in Fig. 7(c) and clearly exhibit the stick-slip behavior observed in macroscopic granular materialsKrim, Yu, and Behringer (2011) as well as in nanoscale single-crystal metal pillarsGreer and Nix (2006). The zero point of the force is determined by recording the positions of the probes in the acoustic field without the raft (for that, the field is turned off after the conclusion of the experiment so that particles fall, and then turned back on and an image is recorded). From the start of the experiment ($t_{0}$), applied force and displacement increase nearly linearly to a peak (near $t_{1}$). This corresponds to the cohesive granular crystal deforming via slight internal rearrangements but without particles slipping past one another. The rapid displacement increase together with a sudden drop in force after the first peak represents a slip event, and thus the onset of significant plastic deformation. A local force minimum occurs after the slip as the crystal recovers. The subsequent rise toward the second peak ($t_{2}$-$t_{3}$) signals the buildup in stress before another slip event ($t_{4}$-$t_{6}$), during which the raft lengthens significantly and is pulled almost completely apart. Attraction between the particles (and between the two large, crystalline portions of the raft) causes the raft to remain contiguous after the event ($t_{7}$). The raft contraction between $t_{6}$ and $t_{7}$ highlights how the acoustic forces can mimic cohesion due to attractions that are longer-ranged than forces based solely on direct contact.
@@ VII Conclusions
The experiments described here were performed in air. Many biological applications of acoustic levitation involve levitating objects in a liquid or gel. The acoustic frequencies used in these systems are typically in the MHz range, with wavelengths 10-100$\times$ shorter than the wavelength in our air-based system. For use in liquids (e.g. in microfluidic systems) the tools described here could be miniaturized and supplemented with more sensitive force sensors based on piezoelectrics, and the procedures could be modified to perform similar tests in different acoustic systems.

### v2

@@ I Introduction
where $r$ is the center-to-center distance and $E_{0}$ represents the acoustic energy density.Silva and Bruus (2014); Lim, VanSaders, and Jaeger (2024) Note that the magnitude of the secondary scattering force is proportional to the product of the two particles’ volumes. While attractive in the nodal plane, the interaction is repulsive normal to that plane, i.e. along the levitation axis. When many objects are levitated, this results in the formation of a close-packed, quasi-two-dimensional membrane or raft.Lim et al. (2022) For more detail on the three-dimensional structure of the secondary acoustic scattering potential, we refer to Lim et al.Lim, VanSaders, and Jaeger (2024)
@@ II.3 Force detection, probe tips, and calibration
A convenient way to calibrate the forces experienced by the probes employs a $\upmu$g-resolution scale (Sartorius CPA26P). During calibration, the plastic handle at the far end of the cantilever wire is attached to a micrometer oriented along the vertical, $z$-direction and lowered so that it barely makes contact with the scale’s platform. The scale’s doors are closed as much as possible, and openings covered with plastic sheeting, to minimize noise from air currents in the room. Incrementally pressing the probe tip onto the scale, mass readings from the scale are recorded and converted to force. The force-deflection data are linear to excellent approximation and fit to a line. The resulting slope is the constant that relates maximum cantilever deflection (or deflection at the probe tip) to measured force. A sample calibration curve is plotted in Fig. 3(c). Alternatively, these probes could be calibrated dynamically by determining the resonance frequency of the cantilevers; we estimate that these sensors have a resonance frequency in the range of 1-10 Hz. The static loading calibration method, however, has the advantage of being independent of any particular model or relationship between applied force and probe deflection.
@@ III Mapping an acoustic cavity
The results of a mapping of this kind are shown in Fig. 4. Here a probe shaped like that pictured in Fig. 4(a) was used, with a 900 $\upmu$m diam. sphere as the tip. The sphere was positioned with its equator in the nodal plane, keeping the rest of the probe tip out of plane. Using a larger probe has the advantage of experiencing a larger primary acoustic force (as primary forces scale with the volume of the scatterer, Wu et al. (2023) see Eq. 1) and also makes it easier to locate in an image, which results in increased relative resolution for determining force values. This probe tip was moved from the center of the cavity along the $x$-axis to a distance greater than $\lambda/2$ and back twice, for four separate trials. For each measurement, we moved the probe into position and then waited for the acoustic pressure control to adjust to changed resonance conditions before recording observations. Videos were recorded with the field on and off for each position. For this and other experiments, image processing to locate probe positions in images was performed with Meta’s Segment Anything model.Kirillov et al. (2023)
@@ IV Measuring the scattering force
Figure 5: Measuring the acoustic scattering force between two objects in the acoustic trap. A T-shaped probe is moved towards a force sensor with a 180 $\upmu$m spherical tip and then pulled away. The spherical probe is on a cantilever to measure forces. The probe positions are tracked to experimentally determine the acoustic scattering force between the probes. (a) Images from the experiment, viewed from below. (a, top) Start of the experiment when the probes are well-separated. (a, bottom) Middle of the experiment when the two probes are almost in contact. Note that the spherical probe has moved slightly to the left as a result of attraction to the larger probe. (b) Schematic of the experimental procedure. The T-probe moves to approach the spherical probe; the spherical probe is attached to a spring-like cantilever to measure the attractive scattering force between the two objects. The eye symbol represents the perspective of the experimental images in panel (a). The long arm of the right probe is not visible in panel (a) because it sits outside the focal plane. (c) Plot of force vs. distance for this experiment (red) and corresponding simulations (blue). Five trials of the experiment were performed, and the data binned by inter-probe distance (error bars: standard error). The scattering force decreases sharply as distance increases. The gray bar marks the noise floor for the experimental force sensor. The force is normalized to allow for comparison between experiment and simulation. The shape of the simulation results is similar to the experimental results. Two triangles are drawn to represent $r^{-3}$ (left) and $r^{-4}$ (right) power laws; these correspond to a simple analytical model of the scenario.
@@ VI Microscale mechanical testing of levitated granular materials
Tension force and displacement data for this experiment are shown in Fig. 7(c) and clearly exhibit the stick-slip behavior observed in macroscopic granular materialsKrim, Yu, and Behringer (2011) as well as in nanoscale single-crystal metal pillars.Greer and Nix (2006) The zero point of the force is determined by recording the positions of the probes in the acoustic field without the raft (for that, the field is turned off after the conclusion of the experiment so that particles fall, and then turned back on and an image is recorded). From the start of the experiment ($t_{0}$), applied force and displacement increase nearly linearly to a peak (near $t_{1}$). This corresponds to the cohesive granular crystal deforming via slight internal rearrangements but without particles slipping past one another. The rapid displacement increase together with a sudden drop in force after the first peak represents a slip event, and thus the onset of significant plastic deformation. A local force minimum occurs after the slip as the crystal recovers. The subsequent rise toward the second peak ($t_{2}$-$t_{3}$) signals the buildup in stress before another slip event ($t_{4}$-$t_{6}$), during which the raft lengthens significantly and is pulled almost completely apart. Attraction between the particles (and between the two large, crystalline portions of the raft) causes the raft to remain contiguous after the event ($t_{7}$). The raft contraction between $t_{6}$ and $t_{7}$ highlights how the acoustic forces can mimic cohesion due to attractions that are longer-ranged than forces based solely on direct contact.
@@ VII Conclusions
Here, we have utilized specialized equipment such as a high speed camera, an optical microphone, and advanced 3D printing techniques. Some of these methods, however, can still be effective without this equipment. For example, quasi-static measurements could be made using cameras with a slower frame rate. Simpler probe structures could be constructed by carefully folding thin wires, though high-resolution 3D printing is required for producing probes of the shapes shown in Figs. 3 and 4.
Extensions of this method could use dynamic excitations near the cantilever’s resonance frequency to measure acoustic forces, similar to frequency modulation atomic force microscopy techniques. Additionally, while the experiments described here were performed in air, many biological applications of acoustic levitation involve levitating objects in a liquid or gel. The acoustic frequencies used in these systems are typically in the MHz range, with wavelengths 10-100$\times$ shorter than the wavelength in our air-based system. For use in liquids (e.g. in microfluidic systems) the tools described here could be miniaturized and supplemented with more sensitive force sensors based on piezoelectrics, and the procedures could be modified to perform similar tests in different acoustic systems.